OpenStudy (laurennarena):

An aircraft travels with the wind for 120 miles in 0.75 of an hour. The return trip is flown against the wind and takes exactly 1 hour. Which system of linear equations represents x, the speed of the plane in miles per hour, and y, the speed of the wind in miles per hour? Recall the formula d = rt.

OpenStudy (laurennarena):

choices: A. 0.75(x + y) = 120 x − y = 120 B. 0.75(x − y)= 120 x + y = 120 C. 120(x + y) = 0.75 120(x − y) = 1 D. 120(x − y) = 0.75 120(x + y) = 1

OpenStudy (adrimit):

the answer is A